OpenStudy (anonymous):

Write this equation in slope-intercept form 3x+6y=15

OpenStudy (anonymous):

Slope Intercept Form: y = mx + b

OpenStudy (anonymous):

y=-1/2x+5/2

OpenStudy (anonymous):

Here is my what I got 6y = -3x + 15 y = (3/6)x - 3 y = (-6/3)x + 3 y = (-3/6)x + 3

OpenStudy (anonymous):

3x+6y=15 subtract 3x from both sides 3x+6y=15 -3x -3x ---- ----- 6y = -3x + 15 Divide by 6 on both sides 6y = -3x + 15 -- ---- --- 6 6 6 \[y = -\frac{ 3 }{ 6 }x + \frac{ 15 }{ 6 }\] Your not done yet, can you simplify and reduce?

OpenStudy (anonymous):

No help me to finish that

OpenStudy (anonymous):

slope intercept in y = mx + b. 3x + 6y = 15. Move terms 6y = -3x + 15. Then divide by 6 to isolate y. so, y = -1/2x + 5/2. hope this helps.